OpenStudy (anonymous):

Help with properties. If n + 4 = 8 + 4 then n = 8 Is this given or substitution or other?

OpenStudy (anonymous):

Hello!

OpenStudy (anonymous):

Is your question of how they got n=8?

OpenStudy (anonymous):

n + 4= 8 + 4 -4 -4 n = 8

OpenStudy (anonymous):

I'm confused as to what property this is?

OpenStudy (anonymous):

Remember that what you do to one side you can do to the other :)

OpenStudy (anonymous):

I'm not sure what property it is...

OpenStudy (anonymous):

Or the name of the property.

OpenStudy (anonymous):

I'm not sure If it's given or if it's substitution

OpenStudy (anonymous):

Well if the problem has this, it is given. You can solve it by substitution or by doing what I did shown above.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i think its most likely given because they give you that n = 8. all you had to do was plug it in the equation :)

OpenStudy (mathstudent55):

Given: n + 4 = 8 + 4 Prove: n = 8 Statements Reasons n + 4 = 8 + 4 Given n = 8 Subtraction property of equality To go from n + 4 = 8 + 4 to n = 8, you subtract 4 from both sides of the equation. You can do that because of the subtraction property of equality.